OpenStudy (inkyvoyd):

Polar functions and derivatives: 1. show that the cardioid r=1+sin(theta) has a horizontal tangent line when r=1/2

OpenStudy (inkyvoyd):

For some reason I don't think I'm doing this right...

OpenStudy (inkyvoyd):

I calculated d(theta)/dr implicitly, getting 1=0+cos(theta)d(theta)/dr sec(theta)=d(theta)/dr Then I tried to express theta in terms of r, so I used the equation and substituted r=1/2 1/2=1+sin(theta) sin(theta)=-1/2 Now what?

OpenStudy (inkyvoyd):

I tried to turn sin(theta)=-1/2 into secant theta, but the answer I'm getting is cos(theta)=pm 2/(sqrt 3)

OpenStudy (inkyvoyd):

@amistre64 , got any ideas what I"m donig wrong?

OpenStudy (asnaseer):

you need to work out dy/dx and set it to zero. see here for a good guide on how to do this: http://tutorial.math.lamar.edu/Classes/CalcII/PolarTangents.aspx

OpenStudy (amistre64):

the puals website is a great resource indeed

OpenStudy (inkyvoyd):

Oh I c, I'm supposed to convert it into a cartesian equation?

OpenStudy (inkyvoyd):

The question is, why can't I work with it while it's in polar form?

OpenStudy (asnaseer):

no

OpenStudy (asnaseer):

no need to convert - look at notes in the link I gave and all will become apparent :)

OpenStudy (anonymous):

i could be wrong, but i think it is \[\frac{dy}{dx}=\frac{\frac{dr}{d\theta}\sin(\theta)+r\cos(\theta)}{\frac{dr}{d\theta}\cos(\theta)-r\sin(\theta)}\]

OpenStudy (asnaseer):

^^ - that is basically the summary result

OpenStudy (asnaseer):

so you just need to calculate dr/d(theta)

OpenStudy (anonymous):

oh look! paul saves the day again. whew.

OpenStudy (asnaseer):

:)

OpenStudy (inkyvoyd):

Oh wow - thanks!

OpenStudy (asnaseer):

plug the result in the equation above, set it to zero and solve

OpenStudy (asnaseer):

there is one slight /twist/ in the solution - hopefully you will be able to see it :)

OpenStudy (anonymous):